Proteins differ in

nitrogen content based on their

AA compositions, but they average about 16 % (w/w)

nitrogen overall. Given this average, calculate the daily diet

ary protein requirement to maintain nitrogen balance

assuming you excrete 140 mg of N/kg body mass each day.
Answer #1

Assuming body weight of a healthy individual =70 kg

N excretion = 140mg/kg so total excretion = 140X70 mg = 980mg

To maintain nitrogen balance, intake and excretion level should be same.

so daily protein requirement is 980 mg.

Now proteins have 16% (w/w) pf Nitrogen meaning 16 mg nitrogen in 100 mg protein

So for 980 mg nitrogen the required protein is = 100X980/16 = 6125 mg = 6.125gm of protein or as per body weight the protein requirement is = 6125/70 mg/kg = 87.5 mg nitrogen /kg of body weight